Cephalexin 25 mg/kg/day PO divided q 8h is ordered for a child with a beta-hemolytic streptococcal infection. The child weighs 18 kg. Your bottle of Cephalexin is labeled 125 mg / 5 mL. How many mL will you give the child every 8 hours? .
Answer #1

Ans) Cephalexin 25 mg/kg/day PO divided q 8h is ordered.

25 × 18 = 450 mg

CCil's weAvailable: 125 mg/ 5 ml

450 mg = x

x = 450 × 5 / 125 = 75 ml

In 3 divided doses, 75/3 = 25 ml every 8th hrly.